Verify that the total differential equation z(z-y)dx+z(x+z)dy+x(x+y)dz=0 is integrable and hence find its integral


1
Expert's answer
2021-02-01T19:07:28-0500

The integrability condition:

"\ud835\udc43 (\n\u2202\ud835\udc44\/\n\u2202\ud835\udc67\n\u2212\n\u2202\ud835\udc45\/\n\u2202\ud835\udc66\n) + \ud835\udc44 (\n\u2202\ud835\udc45\/\n\u2202\ud835\udc65\n\u2212\n\u2202\ud835\udc43\/\n\u2202\ud835\udc67\n) + \ud835\udc45 (\n\u2202\ud835\udc43\/\n\u2202\ud835\udc66\n\u2212\n\u2202\ud835\udc44\/\n\u2202\ud835\udc65\n) = 0"


"z(z-y)(x+2z-x)+z(x+z)(2x+y-2z+y)+x(x+y)(-z-z)="

"=2z^2(z-y)+2z(x+z)(x+y-z)-2zx(x+y)="

"=2z^3-2z^2y+2zx^2+2zxy-2z^2x+2z^2x+2z^2y-2z^3-2zx^2-2xyz=0"

The equation is integrable.


"z(z-y)dx+z(x+z)dy+x(x+y)dz=0"

This equation is homogeneous.

"Px+Qy+Rz=z(z-y)x+z(x+z)y+x(x+y)z="

"=z^2x+z^2y+x^2z+xyz=D\neq0"

Integrating factor:

"\frac{1}{D}=\frac{1}{z^2x+z^2y+x^2z+xyz}"


Then:


"\frac{z(z-y)dx+z(x+z)dy+x(x+y)dz}{z^2x+z^2y+x^2z+xyz}=\frac{z(z-y)dx+z(x+z)dy+x(x+y)dz}{z(x+y)(x+z)}=0"


"\frac{(z-y)dx}{(x+y)(x+z)}+\frac{dy}{x+y}+\frac{xdz}{z(x+z)}=0"


"d(ln(x+y)-ln(x+z))+\frac{dz}{x+z}+\frac{xdz}{z(x+z)}=0"


"d(ln(x+y)-ln(x+z))+\frac{(x+z)dz}{z(x+z)}=0"


"ln(x+y)-ln(x+z)+lnz=lnC"


"\frac{z(x+y)}{x+z}=C"